wiki:UC20200302

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 03/02/2020

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Kal Kadlec (UW), John Bally (CU), Misty Bentz (GSU), Moire Prescott (NMSU), Mukremin Kilic (OU), Eric Bellm (UW), Kevin Schlaufman (JHU), John Wilson (UVa)


User feedback and comments from institutional representatives:

  • Colorado - nothing to report
  • FGCU - no report
  • Georgia State - nothing to report
  • JHU - nothing to report
  • NAPG - no report
  • NMSU - nothing to report
  • Oklahoma - users wanted an update on the status of KOSMOS as they are eager to use it
  • Seattle - nothing to report (by email)
  • UVa - nothing to report
  • Washington - nothing to report
  • Wyoming - no report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 1/30/2020 - 2/26/2020

0) Overview

Numerous winter storms have come through but all have been small and deposited little moisture. Overall there have been no serious telescope or instrument issues. We continue to make substantial progress in planning for a public open house on March 7 with a half-night of eyepiece viewing.

1) Telescope

Nominal operations with minimal problems.

2) Instruments

DIS: The blue camera contamination is still bad and there are no plans to service it. The red camera is at times worse and at times better. There hasn’t been a consistent worsening trend. We continue to monitor this. There were minor issues with the cart mounting again that were addressed.

TripleSpec is operational. A pattern noise was noted at a low level. Ed installed the spare power supply, which improved things quite a bit. The noise level may be present at times but seems to be at reduced and possibly acceptable levels. It does appear more frequent with inversion of the instrument on the NA2 rotator. We are back at the monitoring stage right now.

Agile is operational. There was a report of incorrect image type information in the headers, which we believe is now fixed.

Echelle is operational. The improved inter-order light in the blue has leveled off; red is still good.

NIC-FPS is operational. The reseating of the cards seemed to help with readout issues.

ARCTIC is operational. We had to reset a lost quad again. We also have a header issue that we are working on and can now test better. We have improved the heater to make warmup safer and more efficient. Also with this change the dark current seems to have improved and we are looking at a preferred temperature set point.

ARCSAT: Has been operational. We are in the process of porting some computing tasks over to a new computer to hopefully reduce the amount of computer crashes/hangs in the future. We are also starting to plan for the mounting and first light for the echelle spectrograph.


Additional telescope and instrument discussion:

The new NA2 baffle is installed, having heard no objections to doing so. The IT staff has been working on the header issues for Agile (fixed) and ARCTIC (working on it); these are now at the point where they can be tested. We made changes to the ARCTIC heater current, which looks like it improved the dark current. The goal is to set it to one temperature for year round operations, but the cryotigers seem to be pushed to their limit during the summer months so we're not sure if we can achieve the optimal temperature for ARCTIC during the summer months. For DIS we had to change the cryotiger last Friday. We have older cryotiger compressors for spares; with KOSMOS coming we are reluctant to spend tens of $k if we don’t need to; we are not yet at a critical point but could be within the year. For TripleSpec we made some improvements to the power supply, which reduced the noise levels; we are continuing to monitor this. The pattern noise happens more frequently when the instrument is inverted (but not always), but the noise level seems to be back down to acceptable level (now ~ 7 DN instead of 22). John Wilson suggested that Ed should compare the two power supplies by doing a bench test, and if he finds anything wrong then send the malfunctioning one to Leach for examination/repair. For ARCSAT, the new computer load distribution should improve issues with users’ computers. We are close to having first light on the Shelyak and are looking at possibly commissioning it in April. If users have an interest in using this instrument we would love to have their input on commissioning; please contact Bill Ketzeback (APO) and Derek Buzasi (FGCU) if interested in helping (e.g. by suggesting potential targets and/or helping to develop the commissioning plan). We would like the ARC user community to get involved!


APO and the COVID-19 virus

We have concerns about people who travel and then come to the site as visitors (e.g. for class trips). So far there have been no confirmed cases in NM. The observatory is equipped with masks, gloves, tissues, and wipes and they will be made available to everyone at the site. We have instructed the front office to send a message to everyone who has or will be making housing reservations, requesting that everyone planning to visit the site should follow their institutional guidelines, information from the CDC, and check with the site before coming. If any members of a group planning to visit APO develop symptoms consistent with the virus prior to coming, they should stay home. If any visitors develop symptoms while on site, they should notify Mark immediately so that he can warn the staff.


Updated shutdown schedule

The summer shutdown schedule has changed because of changing availability of the aluminization facility at KPNO: the new summer shutdown dates are July 6 - July 28. This will not result in more time being awarded to institutions at the end of Q2 (the previous dates of the shutdown included the last week of June). We will hold those dates in reserve in the event that KOSMOS becomes ready for commissioning at the end of the quarter.


Q1 3.5m scheduling:

There are 3 half-nights remaining in Q1.


Q1 ARCSAT scheduling:

The ARCSAT time in Q1 has been fully scheduled.


Q2 3.5m scheduling:

The schedule is being built as we speak. Please submit your institutional requests if you have not already done so!


Q2 ARCSAT scheduling:

The call for Q2 ARCSAT proposals will be issued once the Q2 3.5m schedule is completed.


NA2 Baffle Project:

The new NA2 baffle has been installed and tested and it is working well. The initial concerns by some users were that there would be an increase in stray light on NA2 instruments, and that the new baffle would impact the NA2 guider because its FOV is partly blocked by new baffle. ARCTIC would clearly see the greatest benefit because its FOV is no longer vignetted. Preliminary tests show that other instruments with smaller FOVs will not be affected by the presence of the new baffle. The NA2 guider also does not seem to be negatively impacted at all. However, users should know that the configuration of the NA2 port has changed! Flat field frames, for example, could be different before and after the NA2 baffle installation. We welcome any assistance from the user community as people proceed with their data reduction; it would be helpful to receive analyses of the impact of the new baffle on user data. This baffle cannot be changed out quickly, i.e. not during the middle of the night, but if there is a well justified request for putting the older baffle back in we would consider doing that. UC reps should inform their users of the change and request feedback for all NA2 instruments as well as the NA2 guider performance. This can most effectively be done by users who repeat observations from previous sessions to compare data from before and after the new baffle was installed (assuming the same instrument configuration is used). John Bally asked if we could get the observing specialists to take a set of twilight flats in standard filters, since it would be useful to have a standard set of flats for users who don’t have access to twilight time. Mark said he thought we had some, but we would try to collect more during upcoming engineering time.


KOSMOS Update:

The CCD was repaired by e2V - the wire bonds were replaced - and all voltages now read correctly after their in-house testing. It was shipped back to UW last week and we are awaiting its arrival. The TEG is making good progress on the dewar and they think they have identified the cause of leak. They are now working on repairing it.


Status Update on Near-Term Action Items from Strategic Planning Meeting

  • DAWG: No update this month.
  • Getting narrow-band filters for ARCTIC: in progress. Nancy needs to make the purchase.
  • Call for key project proposals: Nancy brought this up as a way to initiate the conversation about how to proceed with this. Several thoughts were shared, including the following:
    • John Bally reiterated that we need to be able to store data on the APO servers for a longer time (e.g. if someone wanted to go back and measure proper motions)
    • in some cases, cross-institution collaborations are already happening naturally (e.g. the UW-OU-CU-NAPG team using ARCTIC + diffuser to observe exoplanet transits)
    • the next step may be to call for white papers to solicit ideas (but we need to announce how much time would be available for such key projects)
    • the observing time would likely be a combination of time that is contributed by member institutions / lease partners as well as some OPEN/DD time
    • for a key project to be successful, it needs to be something that cannot be done using individual PI-led observations; people should be encouraged to think big
    • this discussion may be a bit premature since users may want to assess the performance of KOSMOS before deciding if they want to propose a key project

ACTION ITEMS:

Open action items from previous meetings:

  • Nancy will contact trip leaders bringing student groups to APO in Q1 (-> Q2) to discuss training video ideas. STATUS: OPEN. Nancy will contact Candace and Adam Kowalski from CU, who are planning to conduct a fully online training session, to find out what was missing or most challenging that would have benefited from having some video tutorials. She will communicate that info to Muk (OU), who is bringing a training group to the site in April, so his students can focus on those topics.

New action items from this meeting:

  • UC reps should notify their users of the NA2 baffle switch and request feedback on performance, particularly if any users have before-and-after data that they can use for a quantitative comparison.
  • UC reps should ask if any of their users would like to help with the commissioning of the Shelyak echelle spectrometer on ARCSAT.

All Other Business

Nancy was asked about the echelle advisory committee. Not much has happened yet; she received some suggested names of members but has not yet organized the group. Nancy will contact those people and will distribute the names to 35m-general to see if anyone else wants to join; then she will plan a telecon to kick off the discussion.


Next meeting:

The next Users Committee Zoom meeting will be on Tuesday April 7 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.